# JEL: C81, O32

### v2

@@ 1. Introduction
The idea that patterns like those that motivate this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (a phrase due to Griliches, (1994)) is, similarly, an old one. Early appearances in the literature include Robert K. Merton’s 1935 discussion of the “possibility of a slackening in the rate of technologic advance” (Merton,, 1935), Alfred B. Stafford’s 1952 paper titled “Is the Rate of Invention Declining?” (Stafford,, 1952) and Jacob Schmookler’s various investigations (Schmookler,, 1952, 1954, 1966) into patterns suggestive of diminishing returns to research. Our reconciliation of these patterns hews closely to that of this earlier literature, which viewed the evidence base on productivity decline as inconclusive, given the shortcomings of available measures of innovation (Pakes and Griliches,, 1980; Griliches,, 1990, 1994). [^15] Griliches, (1990) summarizes his version of this conclusion more prosaically: “One can always worry that the world is coming to an end. Someday it undoubtedly will, but it does not look as if the end is already upon us, at least not yet.” This paper differs on one important dimension. We show that advances in data construction technology, if deployed carefully, provide researchers with sufficiently precise measures to draw conclusions.
@@
A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023). We join a small set of papers, e.g., Bartik et al., (2023) and Dell, (2024), that use generative AI as a tool to construct high-quality economic data. Our approach is distinct in this literature, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model. In doing so, we show that it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost. Our use of these technologies is distinct from other applications considered in e.g., Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing.
@@
More generally, we contribute to the large empirical literature that uses text as data (Gentzkow et al.,, 2019). We anticipate that our procedure for the construction of new, precise data on scientific research may be applicable to other contexts where unstructured text is available and manual construction of data is infeasible. We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield data of sufficiently high quality at comparatively low cost.
2. Measuring Research
Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research (e.g., scientists, dollars, research capital) to new, useful information. The canonical Romer, (1990) specification, which motivates a recent empirical investigation in Bloom et al., (2020), is given by
$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=\alpha_{t}S_{t}\leavevmode\nobreak\ .$ (2.1)
Here, the variable $A_{t}$ denotes total factor productivity. The left-hand side of (2.1) captures the rate at which “ideas” are produced. On the right-hand side, the factor $S_{t}$ measures capital devoted to research. The structural parameter $\alpha_{t}$ is interpreted as research productivity—the efficiency with which inputs are transformed into outputs. [^16] See e.g., Romer, (1990), Jones, (1995), and Aghion and Howitt, (1992), among many others.
Although each element of this framework is simple, constructing their empirical analogues is challenging. Decades of scholarship—summarized in Griliches, (1990, 1994) and, more recently, Bryan and Williams, (2021)—have grappled with alternative ways of measuring outputs, $A_{t}$. [^17] Pakes and Griliches, (1980) consider this challenge and employ slightly different notation than we use here. They focus on the quantification of an output variable $K$, which is the “level of economically valuable technological knowledge.” In this paper, our focus is on the comparable challenges associated with quantifying the resources that go into their production, $S_{t}$.
An ideal measure of resources $S_{t}$ should capture the total quantity of economic resources devoted to increasing the stock of useful knowledge. Not all economic activities related to scientific research are equally relevant. Many forms of scientific output have as their goal the synthesis or diffusion of existing information. In medicine, for example, literature reviews and meta-analyses allow physicians to rapidly review the empirical evidence supporting particular treatment options. [^18] For one discussion of the appropriate interpretation of meta-analyses in clinical practice, see Thompson and Pocock, (1991). Medical guidelines, produced by entities like the American Heart Association (e.g., Smith Jr. et al.,, 2001), consolidate these findings into consensus documents that may be viewed as defining the “standard of care” (Moffett and Moore,, 2011). These types of literature syntheses are, anecdotally, of increasing importance, as health technologies previously available only in a small number of countries are made available to new patient populations (Ioannidis et al.,, 2013; Cockburn et al.,, 2016).
It is important to be clear about the distinction we are making between these forms of research. This paper, and productivity calculations more generally, are interested in those forms of research that generate new information. These are the elements of the quantity $S_{t}$. From a welfare perspective, there are many other types of investment that are important, including these types of literature reviews that aid in the diffusion of information. These other investments may enable the “resources related by productivity growth . . . [to] find adequate employment,” thus increasing aggregate welfare, but are not themselves contributing to productivity growth (Griliches,, 1994).
Even if, conceptually, it is clear that certain economic activities are more relevant to the quantity $S_{t}$ than others, available indicators of research—such as scientific publications and patents—may make it challenging to accommodate this heterogeneity empirically. This task is somewhat simpler in industries often studied by economists (see, e.g., the discussion of manufacturing firms in Syverson,, 2011), where dollar-denominated expenditures and price indices are available. In many settings central to economic growth, however, research occurs in university-based laboratories that do not maintain detailed financial ledgers, [^19] The UMETRICS data are one exception. See Babina et al., (2023) for an example of a study employing financial data from a university research setting. in countries where high-quality data may not be collected or made available, or in industries where confidentiality restrictions limit the public release of dollar-denominated expenditure records. Consequently, measures of research must rely on counts of researchers (who differ in skill), publications (which differ in relevance and resource-intensity), and administrative databases (which differ in completeness).
The objective of this paper is to address these conceptual and empirical challenges in the context of medical research. Measuring the outputs of the production function (2.1) is, in some sense, more straightforward in medicine than in other sectors. Medical innovation appears coarsely in records maintained by regulators (e.g., counts of “new molecular entities” approved by the U.S. Food and Drug Administration) and, more directly, in measures of life expectancy and life-years saved from disease. [^20] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases. In a series of case studies, they argue that the benefits of technological progress exceed their costs. See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin. As this paper focuses on quantification of inputs $S_{t}$, we note only that it is a well-established empirical regularity that improvements in life expectancy, on many margins, appear to be “stagnating” (Mehta et al.,, 2020; Ho and Hendi,, 2018). In drawing conclusions about research productivity, we will follow Bloom et al., (2020) and assume that the left-hand side of (2.1) is constant. [^21] It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
Our task, then, is to identify a measure for inputs $S_{t}$ that best explains potential shifts in this output measure. Institutional features of medicine are helpful in doing so: medical decision-makers enforce a clear hierarchy of statistical evidence (Jones and Podolsky,, 2015). As guidelines to medical journal authors and editors from the American Medical Association state, “[p]rospective studies generally yield more reliable conclusions than retrospective studies. Cause and effect cannot be established by observational studies; consequently, study findings from observational studies should not be presented using causal language.” [^22] In 2024, the Journal of the American Medical Association—the “trunk journal” (see Section 4) of the American Medical Association and among the highest ranked journals in medicine—adopted a framework to allow studies other than randomized experiments to use this type of “causal” language. In doing so, editors of JAMA noted that “well-conducted randomized clinical trials remain the preferred approach” for the production of this type of evidence (Dahabreh and Bibbins-Domingo,, 2024). In decisions about the approval of new medicines (Food and Drug Administration,, 2024), the terms of insurance coverage (Choudhry,, 2017), and the appropriateness of a particular prescription (Martí-Carvajal,, 2020; Alsan et al.,, 2024), decision-makers defer to evidence generated via clinical trial. Other forms of medical evidence, generated through observational studies and surveys of the literature, receive less weight. Although detailed cost information on medical research is not available, available evidence suggests that these other forms of medical research are lower cost [^23] Recent estimates suggest that the average clinical trial may cost upwards of \$20 million dollars (Martin et al.,, 2017). Available estimates suggest that the cost of completing a systematic review or meta-analysis can reach \$150,000, where the bulk of the cost stems from the quantity of labor required to identify trials in publication data (Michelson and Reuter,, 2019). and involve fewer regulatory hurdles (Congressional Budget Office,, 2021). [^24] For example, while the approval of an institutional review board is necessary for any randomized clinical trial to proceed, many observational studies are exempt from this type of review. See guidance to researchers at Stanford University here: https://stanfordmedicine.box.com/shared/static/ase46ixmatdeahj7qdfgp4ydmy5qc3ei.pdf.
Thus, we treat clinical trials as the appropriate measure of research inputs $S_{t}$. Clinical trials are discrete, regulated experiments. Research proceeds according to a pre-defined experimental protocol under the supervision of a regulator and across various phases. Trial phases differ in their scale (number of participants and duration) and, thus, in their cost (Food and Drug Administration,, 2024). Standardized measures (e.g., statistical objects like p-values) are recorded at various intervals. Trials are sufficiently expensive (Congressional Budget Office,, 2021) as to reflect “real” and substantial investment, and decisions to progress a study across phases (e.g., to move from a small Phase II study to a large Phase III study) capture decisions to pursue continued, directed research in an area of study (for an application of this insight, see Cunningham et al.,, 2021).
As we approach this exercise with the ultimate goal of making progress on the question of research productivity, it is useful to note an additional feature about clinical trials, which suggests that sharp increases in quantities year-on-year (as documented in Bloom et al., (2020)) are unlikely. Not only are clinical trials sufficiently expensive as to require access to a source of financing, but trials also require an often scarce “input:” eligible patients. Malani and Philipson, (2019) collect empirical evidence of this type of “subject market effect,” and Alsan et al., (2024) collect anecdotal evidence on the difficulties of patient recruitment. [^25] The popular press has described this as a conundrum for cancer research: “too many drug trials, too few patients” (Kolata,, 2017).
@@ 3. Identifying Clinical Trials
We construct a new census of clinical trials disclosed in scientific publications. We begin, in Section 3.1, by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical research. In Sections 3.2 and 3.3 we propose a method, based on fine-tuning a large language model, for identifying relevant records at scale and with high fidelity. The performance of the method is assessed in Section 3.4.
@@ 3.1. Existing Approaches
Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent quantity measures for medical research in aggregate.
3.1.1. Dollar Denominated Data
Although disaggregated, dollar-denominated trial costs are a natural measure, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently (for a discussion, see Morten et al.,, 2024).
There are, to our knowledge, three such sources of data that have been used in research. DiMasi et al., (2003, 2016) produce one widely cited (and contested, see Light and Warburton, (2005)) set of estimates of drug research and development costs, based on confidential data obtained from 10 firms. Myers and Pauly, (2019) and Scannell et al., (2012) collect aggregate estimates of annual research and development expenditures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry. Both sets of estimates are often characterized as an upper bound (Morten et al.,, 2024) and cannot be produced independently by external researchers. Wouters et al., (2020) generates estimates of research and development costs, drawing on the information available in financial filings collected by regulators in the United States. Although these estimates are important and informative in a context that lacks precise measures of costs, expenditures can be defined in this way for only roughly 18 percent of the new drugs approved by the FDA between 2009 and 2018.
These estimates, then, are not sufficient to answer the question this paper sets out to address. Even if dollar-denominated measures were available, it is not obvious how these measures should be scaled to allow for comparison across time periods. See Berndt and Cockburn, (2013) for a discussion of the difficulties associated with constructing a price index for clinical trial research. [^26] Berndt and Cockburn, (2013) construct a preliminary price index using data on trial budgets from a private firm. This firm no longer discloses data to academic researchers.
@@ 3.1.2. Trial Registries and Proprietary Databases
In practice, there is widespread noncompliance with these regulations and requirements. Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017). Noncompliance with federal reporting requirements is similarly widespread. In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline. [^27] The FDA is empowered, by statute, to impose fines of up \$10,000 for each day that required disclosures are late. To date, no fines have been collected, but the agency now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021). Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
@@ 3.1.3. Publication Data
Publication data provide a natural alternative. [^30] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time. Neither is likely to affect the interpretation of trends in our data. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^31] See Section A.1 for further details on on our treatment of the PubMed data. It is reasonable to expect that the corpus of scientific publications contains a census of research. The challenge is in extracting measures of research from this unstructured text. Of these $34$ million records, which are clinical trials?
@@
Figure 3.1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition. That few of these series suggest “smooth” trends in the quantity of trials indicates, also, the extent to which the quality of data can affect these estimates. See Section A.2 for further details on the construction of these series.
Given these difficulties, we describe a procedure, below, to recover a sample of clinical trials from this corpus of text in a manner that is consistent and transparent.
@@ 3.2. Objective
3.1 embeds several restrictions, intended to yield a measure of new investment in information relevant to human-facing medical care. We view this set of studies as the set most likely to generate information relevant to human mortality. Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded. We take the “potential universe” plotted in Figure 3.1 as our baseline sample of potential clinical trials.
We further restrict attention to the 1,821,429 records published in or after 2010, as there is a trade-off between data coverage and quality in this setting. In the 1990s and 2000s, medical experts created a standardized system for clinical trial reporting in scientific publications. By the time of the most recent Consolidated Standards of Reporting Trials (CONSORT) in 2010, nearly all medical journals required authors to comply with a standardized structure for all clinical trial publications (Schulz et al.,, 2010). As a result, most trial publications have a distinctive format and include a pre-defined set of elements. Although our procedure can be applied to any type of scientific publication text, the standardization required by the CONSORT Statement changes the relationship between clinical trial research and publications in the early 2000s. To answer the research question that motivates this paper—is the quantity of medical research rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
@@
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 3.2. Further details on the construction of this figure are given in Section C.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
@@
Notes: Figure 3.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1. The hyper-parameters of each model are determined with 10-fold cross-validation. The models are trained in the training and validation data set. Error rates are estimated in the testing data set. In addition, we display the performance of our, ensemble, fine-tuned large language model in purple. This is the same curve labelled “Ensemble” in Figure 3.3.
3.3. Prompt Design and Fine-Tuning
We construct a large language model optimized for our task. This is accomplished in three steps. First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4. A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4. The extraction of GPT-3.5 labels is substantially faster and cheaper. Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models. Third, the noisy labels are used to train an off-the-shelf large language model. The resulting model is then used to identify clinical trials in our baseline sample. This process of constructing specialized large language models is referred to as “fine-tuning” (see e.g., Taori et al.,, 2023, for a prominent application of this idea).

### v3

@@ 1.1. Related Literature
The idea that patterns like those considered in this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (Griliches,, 1994) is old and closely tied to the development of methods for the measurement of innovation. Early appearances of this concern in the literature include Robert K. Merton’s 1935 discussion of the “possibility of a slackening in the rate of technologic advance” (Merton,, 1935), Alfred B. Stafford’s 1952 paper titled “Is the Rate of Invention Declining?” (Stafford,, 1952) and Jacob Schmookler’s various investigations (Schmookler,, 1952, 1954, 1966) into patterns suggestive of diminishing returns to research. Our treatment of these patterns hews closely to that of earlier literatures, which view the evidence base on productivity decline as inconclusive, given the shortcomings of available data (e.g., Pakes and Griliches,, 1980; Griliches,, 1990, 1994; Cockburn,, 2006). [^10] Griliches, (1990) summarizes his version of the conclusion that any finding of diminishing returns to innovative efforts are limited by the availability of high-quality data more prosaically: “One can always worry that the world is coming to an end. Someday it undoubtedly will, but it does not look as if the end is already upon us, at least not yet.” This paper differs on one important dimension. We show that advances in data construction technology, if deployed carefully, can allow researchers to resolve some long-standing data construction challenges.
@@
A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023). We join a small set of papers—notably, Bartik et al., (2023) and Dell, (2024)—that use generative AI as a tool to construct high-quality economic data. Our approach is distinct, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model. In doing so, we show that, with model distillation, it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost. Prominent examples of model distillation in the computer science literature focus on the construction of chat bots by fine tuning open source models with data extracted from proprietary models (Taori et al.,, 2023; Chiang et al.,, 2023; Xu et al.,, 2023). Other papers have fine-tuned models for simple completion tasks like arithmetic (Liu and Low,, 2023) or querying an API (Patil et al.,, 2023). We are distinguished from this literature by considering a problem with, arguably, a higher-degree of complexity: the construction of data of sufficient quality to be of substantive interest to social science.
Our use of large language models is also distinct from other applications considered in papers such as Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing. It is worth emphasizing that, in this paper, we develop a procedure that is applicable to a binary classification problem, albeit one that is highly complex. The viability of these methods, and the extent of required modifications from the process outlined here, for application in more complicated classification problems of interest to economists, including those outlined in Dell, (2024), is a valuable direction for further research.
@@
More generally, we contribute to the large empirical literature that uses text as data (Gentzkow et al.,, 2019). We anticipate that our procedure for the construction of new, precise data on scientific research may be applicable to other contexts where unstructured text is available and manual construction of data is infeasible. We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield high-quality data at comparatively low cost.
@@ 2. Identifying Clinical Trials
We construct a new census of clinical trials disclosed in scientific publications. We begin, in Section 2.1 and Section 2.2, by providing context on the objectives of our data construction exercise and by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical innovation. In Sections 2.3 and 2.4 we propose a method, based on distilling data from a large language model, for identifying relevant records at scale and with high fidelity. The performance of the method is assessed in Section 2.5.
2.1. Context
Brief institutional context on the existing evidence related to pharmaceutical research productivity is helpful both in motivating our data construction and in interpreting our findings.
Declining pharmaceutical research productivity has been a topic of policy concern since the late 1990s. In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” R&D, in this setting, refers to a particular type of biomedical research: investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients. That inputs per new drug appear to be rising sharply over time is viewed, generally, as an established fact in the empirical productivity literature (Bloom et al.,, 2020; Goldin et al.,, 2024).
Measures of productivity in this setting are somewhat crude. The measure of outputs in the most influential articulation of this phenomenon, by Scannell et al., (2012), is the number of “new molecular entities”—innovative drugs—approved by the U.S. Food and Drug Administration. [^11] Technically, a new molecular entity is a drug that contains no previously approved active moiety. In practice, this means that every active element of the drug (e.g., acetaminophen and caffeine in the commonly used over-the-counter migraine medicine Excedrin) must be novel. As Bloom et al., (2020) note, that this count has been stagnant or declining is viewed as an established fact. [^12] In data on new molecular entities approved by the FDA between 2005 and 2021, we observe a small increase in the number of approvals, largely driven by an uptick in approvals in 2020 and 2021. In earlier periods of data, we confirm that there is no discernible trend in new molecular entity approvals.
Measures of inputs are, typically, some estimate of R&D expenditures. The simplest construction of a measure of R&D expenditure scales the (inflation-adjusted) cost of a clinical trial by the number of clinical trials. Recent evidence on these objects comes from Bloom et al., (2020) and Sertkaya et al., (2024). Bloom et al., (2020) find a sharp increase in the number of clinical trials, as measured in publication data over time. Using granular data on clinical trial site contracts and internal records of the U.S. Food and Drug Administration, Sertkaya et al., (2024) find that the cost of clinical trials and the cost of drug development was roughly constant between 2000 and 2018.
Arguments of declining productivity in the pharmaceutical industry, then, depend substantially on the finding that the quantity of clinical trials is rising.
@@ 2.2. Existing Approaches
Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent aggregate quantity measures.
@@ 2.2.1. Trial Registries and Proprietary Databases
In practice, there is widespread noncompliance with these regulations and requirements. Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017). Noncompliance with federal reporting requirements is similarly widespread. In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline. [^13] Note that noncompliance in this industry persists despite the existence of high statutory penalties. Sponsors can be fined up to \$10,000 per day for failing to register studies and disclose results. See the FDA Amendments Act of 2007 (FDAAA) and its Final Rule issued in 2021, at 42 U.S.C. § 282(j)(5)(C)(i). To date, no fines have been collected, but the FDA now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021). Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
@@ 2.2.2. Publication Data
Publication data provide a natural alternative. [^16] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time. Neither is likely to affect the interpretation of trends in our data. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^17] See Section A.1 for further details on on our treatment of the PubMed data. The challenge is to classify the specific forms of research captured by this corpus. Of these $34$ million records, which are clinical trials?
@@
Figure 2.1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition. See Section A.2 for further details on the construction of these series. Given these difficulties, we describe a procedure, below, to recover a sample of clinical trials from this corpus of text in a manner that is consistent and transparent.
@@ 2.3. Objective
2.1 embeds several restrictions, intended to yield precise measures of clinical trials relevant to the approval of new pharmaceuticals for use in human patients. Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded. We take the “potential universe” plotted in Figure 2.1 as our baseline sample of potential clinical trials.
We further restrict attention to the 1,821,429 records published in or after 2010, as there is a trade-off between data coverage and quality in this setting. In the 1990s and 2000s, medical experts created a standardized system for clinical trial reporting in scientific publications. By the time of the most recent Consolidated Standards of Reporting Trials (CONSORT) in 2010, nearly all medical journals required authors to comply with a standardized structure for all clinical trial publications (Schulz et al.,, 2010). As a result, most trial publications have a distinctive format and include a pre-defined set of elements. Although our procedure can be applied to any type of scientific publication text, the standardization required by the CONSORT Statement changes the relationship between clinical trial research and publications in the early 2000s. To answer the research question at the center of our empirical analysis—is the quantity of clinical trials rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
@@
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 2.2. Further details on the construction of this figure are given in Section C.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. Performance is measured in the testing sample split. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
@@
Notes: Figure 2.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 2.1. The hyper-parameters of each model are determined with 10-fold cross-validation. The models are trained in the training and validation data set. Error rates are estimated in the testing data set. In addition, we display the performance of our, ensemble, fine-tuned large language model in purple. This is the same curve labelled “Ensemble” in Figure 2.4.
2.4. Model Distillation
We construct a large language model optimized for our task. This is accomplished in three steps. First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4. A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4. The extraction of GPT-3.5 labels is substantially faster and cheaper. Second, we extract noisy labels for a moderate number of publications in our sample by querying the proprietary models. Third, the noisy labels are used to train an off-the-shelf large language model. The resulting model is then used to identify clinical trials in our baseline sample. This process for building specialized large language models is referred to as model distillation (Xu et al.,, 2024).